In 4 years time, Khairu's mother will be three times as old as khairu. six years ago, his mother was seven times as old as him. Find their present ages.
Answer by lwsshak3(11628)

In 4 years time, Khairu's mother will be three times as old as khairu. six years ago, his mother was seven times as old as him. Find their present ages.
***
6 yrs ago:
x=khairu's age
7x=mother's age
..
At present:
x+6=khairu's age
7x+6=mother's age
..
In 4 yrs time:
x+10=khairu's age
7x+10=mother's age
7x+10=3(x+10)
7x+10=3x+30
4x=20
x=5
x+6=11
7x+6=41
khairu's age=11
mother's age=41